OpenStudy (anonymous):

lim x->0 Sin t / 1 + cos t = ?

OpenStudy (anonymous):

There's nothing tricky there, it's a straightforward limit.

OpenStudy (inkyvoyd):

evaluation by substitution...

OpenStudy (inkyvoyd):

Unless you are saying its lim x->0 sin t/(1+ cos t)

OpenStudy (anonymous):

That's how I interpreted it, but it's still straightforward.

OpenStudy (inkyvoyd):

yeah good point there heh

OpenStudy (anonymous):

and I also of course assume it's the limit as t->0

OpenStudy (inkyvoyd):

yeah

OpenStudy (anonymous):

sorry first time using this site, i am not sure i understand what you are saying

OpenStudy (inkyvoyd):

put in values and see what you get.

OpenStudy (anonymous):

but you cant do that because if you plug the limit in you get 0

OpenStudy (inkyvoyd):

Sin t / 1 + cos t t=0 0/(1+1)=0/2=0

OpenStudy (inkyvoyd):

Sin t / 1 + cos t t=0 0/1+cos(0)=1

OpenStudy (anonymous):

right, so is the limit 0?

OpenStudy (inkyvoyd):

I dunno, you didn't use parenthesis to indicate what you were trying to say and usually people get order of operations skewed

OpenStudy (anonymous):

I know how you got to 0/2

OpenStudy (anonymous):

its (sin t)/ 2+ cos t)

OpenStudy (inkyvoyd):

... you're missing a left parenthesis somewhere

OpenStudy (anonymous):

2+ (cos t) is the denomiator

OpenStudy (anonymous):

sorry denominator*

OpenStudy (inkyvoyd):

give me the entire expression with correct usage of parenthesis please -the answer varies greatly otherwise

OpenStudy (anonymous):

lim x->0 (sin t)/ 1 + (cos t)

OpenStudy (inkyvoyd):

=0/1+1=1

OpenStudy (anonymous):

0/2 is 0 not one

OpenStudy (anonymous):

Based on this conversation I assume you're looking for \[\lim_{t\rightarrow 0} \frac{\sin(t)}{1 + \cos(t)} = 0\] The reason there is confusion, dom, is because when you write that you need to put the entire denominator in parentheses lim t->0 sin(t) / (1 + cos(t) )

OpenStudy (anonymous):

Is the answer 0?>

OpenStudy (anonymous):

Yes, the answer is zero.

OpenStudy (anonymous):

TY

OpenStudy (anonymous):

0